# CruxEval-X: A Benchmark for Multilingual Code Reasoning, Understanding and Execution

### v1

@@ Abstract
Code benchmarks such as HumanEval are widely adopted to evaluate Large Language Models’ (LLMs) coding capabilities. However, there is an unignorable programming language bias in existing code benchmarks – over 95% code generation benchmarks are dominated by Python, leaving the LLMs’ capabilities in other programming languages such as Java and C/C++ unknown. Moreover, coding task bias is also crucial. Most benchmarks focus on code generation capability, while benchmarks for code reasoning (given input, reasoning output; and given output, reasoning input), an essential coding capability, are insufficient. Yet, constructing multi-lingual benchmarks can be expensive and labor-intensive, and codes in contest websites such as Leetcode suffer from data contamination during training. To fill this gap, we propose CruxEval-X, a multi-lingual code reasoning benchmark that contains 19 programming languages. It comprises at least 600 subjects for each language, along with 19K content-consistent tests in total. In particular, the construction pipeline of CruxEval-X works in a fully automated and test-guided manner, which iteratively generates and repairs based on execution feedback. Also, to cross language barriers (e.g., dynamic/static type systems in Python/C++), we formulated various transition rules between language pairs to facilitate translation. Our intensive evaluation of 24 representative LLMs reveals the correlation between language pairs. For example, TypeScript and JavaScript show a significant positive correlation, while Racket has less correlation with other languages. More interestingly, even a model trained solely on Python can achieve at most 34.4% Pass@1 in other languages, revealing the cross-language generalization of LLMs. The leaderboard is available at https://cruxeval-x.github.io/leaderboard.html.
Introduction
@@
Large language models (LLMs) have shown advanced proficiency in various domains, including code generation (Liu et al. 2024; Du et al. 2024), defect detection (Yang et al. 2024b) and program repair (Xia and Zhang 2023; Zhong, Wang, and Shang 2024; Hu et al. 2024). Benchmarks such as HumanEval (Chen et al. 2021) and SWE-bench (Jimenez et al. 2023) were introduced to measure LLMs’ capabilities, providing insights into LLMs’ strengths and weaknesses.
@@
Recent studies (Cao et al. 2024a; Chai et al. 2024; Chen et al. 2024) have spotted two significant biases in current benchmarks. First, Programming language bias. As pointed out by prior studies (Cao et al. 2024a; Chai et al. 2024; Chen et al. 2021; Austin et al. 2021), Python dominates code generation benchmarks with over 95% involvement. Other programming languages (PLs) such as Java and C/C++, despite their popularity and availability, gain less exploration. Second, Coding task bias. Most coding benchmarks focus on code generation tasks (i.e., giving descriptions in natural language and generating the program, as shown in Figure 1 (A)), while code reasoning (i.e., given the program, reasoning the input or output of the program, as shown in Figure 1 (B - D)), as an essential coding capability of LLMs, is seldom evaluated (Chen et al. 2024). A recent work introduced a code reasoning benchmark (Gu et al. 2024; Chen et al. 2024), while it is only in Python. Figure 1 (C - D) shows that simply changing PLs from Python to C++ can turn a correct reasoning into an incorrect one.
@@
However, constructing multi-lingual benchmarks is not a trivial task. First, human annotation can be expensive. As reported by recent work (Chai et al. 2024), they spent a total of \$12,000 US dollars for human annotators, providing the working environment, free meals, souvenirs, and free GPT-4 interface usage to construct their multi-lingual benchmark. Second, automated translation does not perform well. The latest studies (Yin et al. 2024) show that even the best LLM (i.e., ChatGPT) can only achieve an average of 64% success translation rate, which is far from practice. Rule-based translation (Cassano et al. 2023; Ling et al. 2022) usually suffers from generalizability issues, making them limited in handling prescribed code structures. Additionally, multi-lingual solutions from contest websites such as LeetCode and Codeforces were included in most LLMs training sources, thus suffering from data contamination issue (Cao et al. 2024b).
@@
To fill the research gaps, we introduce CruxEval-X, a multi-lingual code reasoning benchmark that contains 19 popular PLs, including C++, Rust, Java, etc. , expanded from CruxEval (Gu et al. 2024), a code reasoning benchmark written in Python. For each PL in CruxEval-X, there are at least 600+ functions. In total, there are 12,660 subjects along with 19K test cases for input/output reasoning.
@@
Noteworthy that the pipeline of constructing CruxEval-X works in a fully automated manner. It first translates the test cases by transition rules adapted from prior work (Cassano et al. 2023), then iterates the generation-and-repair process intensively. In particular, the transition rules are formulated to cross the language barriers. For example, Python employs a dynamically typed system where types are determined at runtime, whereas C++ uses a statically typed system requiring explicit type declarations at compile time. The rules facilitate the translation of the test cases. Additionally, inspired by prior work (Yin et al. 2024; Rozière et al. 2022), we employ a test-guided manner (Rozière et al. 2022) to generate the translation and iteratively repair the generated code using execution feedback (e.g., compilation error, runtime error) (Yin et al. 2024).
@@
Through intensive experiments on 24 mainstream LLMs, we observe several interesting findings. First, in multiple PLs, the input reasoning and output reasoning capabilities of LLMs are comparable. Also, there is a noticeable correlation between certain PLs (e.g., JavaScript and TypeScript show a positive correlation, while Racket consistently yields the worst results). More interestingly, we observe that even if a model is only trained on Python (e.g., phi-1 and phi-1.5), it still can reach a 16% $\sim$ 26% output reasoning success rate in other PLs, compared with 25.6% in Python. The finding indicates the cross-language generalization of LLMs.
@@
The contributions can be summarized as follows: (1) We introduce CruxEval-X, a multi-lingual code reasoning benchmark that contains 19 popular PLs. (2) We introduce an automated code translation pipeline that adopts a test-guided and iterative generate-and-repair practice. (3) We evaluate 20+ LLMs against CruxEval-X and yield inspiring findings.
@@
Benchmark Construction
In this section, we detailed the construction process of CruxEval-X in Figure 2. It can be divided into three main steps. First, we translate the function signature via mapping variable type annotations (Step 1 in Figure 2). Then we employ a rule-based approach to translate Python test cases into other PLs (Step 2 in Figure 2). Finally, we integrate multiple LLMs to translate the code by iterating the generation-and-repair process (Step 3 in Figure 2).
Step I. Function Signature Translation
@@
Then, we adopt the rules as prior work (Cassano et al. 2023) to map the types from Python to other PLs. In particular, we identify the data types in the annotated Python signature (e.g., parameter types, return types), mapping the types from Python to other PLs according to the rules, then structuring the signature in the corresponding PLs. Take the example in Figure 2 Step I, the Python signature def f(s1:str, s2:str) -> str is translated into that in C (std::string f(std::string s1, std::string s2). After translating the tests, all 800 subjects in Python can be translated, as shown in Table 1.
Step II. Test Suites Translation
@@
We employ a test-guided approach to ensure the correctness of the translation results, which necessitates test cases in various PLs. Prior works (Athiwaratkun et al. 2022; Cassano et al. 2023) provided various rules for mapping Python test cases to other PLs. We adopt the mapping rules from MutiPL-E (Cassano et al. 2023) to assist the transition of our test suites.
@@
However, their rules have limited support for type handling (e.g., they cannot handle a list with hybrid types), Thus, to maximize the success rate, we made two improvements to enhance the rules. First, we enhance handling structured types such as List and Dict. For example, when handling C#, we add an equality function to check whether two Dict types are equal. Second, when dealing with variables that have complex types that are not as well-supported in some other PLs, we transform these variables into more generic types without significantly altering the original function’s functionality. For example, we change type List[Union(str, int)] into List(str) if the function keeps the same functionality. A small portion of the data that cannot be converted is discarded. The result of Step II is shown in Table 1. Further details can be found in the Appendix.
Step III. Iterative Generation & Repair
After the test, dependencies, and signatures are properly transited, we then use multiple LLMs to iteratively translate the original Python reference code into corresponding PLs. In particular, each LLM undergoes two substeps: generation and repair. These substeps are based on the results of the previous LLM. Once a problem successfully passes the test cases, it will not appear in the next round of iteration. Below, we will elaborate on these two substeps in detail.
Generation.
Relying solely on a single LLM’s limited number of generations is unlikely to achieve high accuracy in translation tasks (Yin et al. 2024). Therefore, we propose a multi-round generation method, which involves interaction with the testing environment to determine whether to proceed with the next round of iteration.
@@
Let $A_{0}$ represent the number of translated codes that pass the tests and the number of total problems is $U$. For all questions that are not correctly answered, we used LLM $M$ to generate results through multiple rounds. We denote the number of correct results in the $i$-th round as $A_{i}$, with the maximum number of generation rounds set as $N$. For $A_{i}$, if the increase in the number of correct questions compared to the results $k$ rounds before $A_{i-k}$ is less than the threshold $\delta$, the process will stop early. Here, $k$ is a fixed constant, and the formula is as follows:
@@
We obtain the $i$-th round code result $O_{i}$ from the code generation distribution $P(O_{i}\ |\ U-A_{i-1};M)$. Then we calculate the correct results using $Correct(.)$ as the function.
To fully leverage the strengths of various LLMs, we select the closed-source LLM GPT3.5-Turbo and the open-source LLM DeepseekCoder-33B-Instruct for iterative translation tasks. We initially use GPT3.5-Turbo for preliminary generation. Given the higher usage cost of closed-source LLMs, we set $N$ to $5$, $k$ to $5$, $\delta A$ to 00, and the temperature to 0.2. The result is shown in Table 1 under the column “w/o Iter”. We also use this as a baseline for the effectiveness of single LLM’s generation within a limited number of iterations to compare with the final results of our pipeline. Subsequently, we employ Deepseekcoder-33B-Instruct on top of the foundation laid by GPT3.5-Turbo to conduct further generation. We set $N$ to $50$, $k$ to $5$, $\delta A$ to 00, and the temperature to 0.8.
Repair
@@
Simply generating code will still result in many errors the LLM cannot solve. Therefore, after the generation step of each LLM, we provide them error messages for error correction. We observed that the cost of multiple iterative error corrections is high and the benefits are relatively low (Chen et al. 2023), so we only perform error correction once after each of the two LLMs.
@@
After the generation of GPT3.5-Turbo, we directly provide the LLM with the erroneous code along with the error messages for correction. After the Iterating Generation of DeepseekCoder-33B-Instruct, since the untranslated code can produce numerous incorrect code snippets after multiple rounds of iteration, which may contain the same errors, we first use simhash to deduplicate the erroneous code and then proceed with error correction on the deduplicated code. Error correction in different phases uses the LLM employed in that specific phase, with a temperature setting of 0.
Multiturn Repair based on overlap
@@
After completing the steps above, we first calculate the intersection of correctly answered questions across different PLs. To our surprise, there were only 333 questions that all PLs answered correctly. However, 563 problems have been successfully translated correctly by at least 16 PLs. Upon analyzing the questions our LLM failed to solve, we find that each PL has its difficulties in translating from python. For example, in Julia, the index for arrays and other collection types starts from 1, which differs from Python. The details of the difficulties can be found in Appendix.
@@
Based on these observations, we conduct final generation and iterative error correction on the questions that are correctly translated by more than 15 PLs. Due to the small number of questions requiring translation, we utilized GPT-4o for generation and error correction. We set the temperature to 0, generate once, and correct errors three times. During the generation process, we provide the LLM with three corresponding typical examples based on the difficulties we find. The overlap is increased to 462 after repair of GPT-4o.
@@
We manually modified 38 questions that GPT-4o almost got right, expanding our dataset to 500 entries. We determined that 500 entries are sufficient to distinguish the effectiveness of the LLMs and most of the questions that failed to be translated cannot be expressed well in other PLs. Therefore, we use these 500 entries as our CruxEval-X benchmark. The final result of our pipeline is shown in Table 1 under the column “w/ Iter”. The prompt of each step can be found in Appendix.
@@
Experiments
Experiment Setup

### v2

@@ Abstract
Code benchmarks such as HumanEval are widely adopted to evaluate Large Language Models’ (LLMs) coding capabilities. However, there is an unignorable programming language bias in existing code benchmarks – over 95% code generation benchmarks are dominated by Python, leaving the LLMs’ capabilities in other programming languages such as Java and C/C++ unknown. Moreover, coding task bias is also crucial. Most benchmarks focus on code generation capability, while benchmarks for code reasoning (given input, reasoning output; and given output, reasoning input), an essential coding capability, are insufficient. Yet, constructing multi-lingual benchmarks can be expensive and labor-intensive, and codes in contest websites such as Leetcode suffer from data contamination during training. To fill this gap, we propose CruxEval-X, a multi-lingual code reasoning benchmark that contains 19 programming languages. It comprises at least 600 subjects for each language, along with 19K content-consistent tests in total. In particular, the construction pipeline of CruxEval-X works in a fully automated and test-guided manner, which iteratively generates and repairs based on execution feedback. Also, to cross language barriers (e.g., dynamic/static type systems in Python/C++), we formulated various transition rules between language pairs to facilitate translation. Our extensive evaluation of 24 representative LLMs reveals the correlation between language pairs. For example, TypeScript and JavaScript show a significant positive correlation, while Racket has less correlation with other languages. More interestingly, even a model trained solely on Python can achieve at most 34.4% Pass@1 in other languages, revealing the cross-language generalization of LLMs.
CruxEval-X: A Benchmark for Multilingual Code
Reasoning, Understanding and Execution
Ruiyang Xu1,2,*, Jialun Cao3,*, Yaojie Lu1, Ming Wen4, Hongyu Lin1, Xianpei Han1, Ben He1,2, Shing-Chi Cheung3, Le Sun1 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3The Hong Kong University of Science and Technology 4Huazhong University of Science and Technology {xuruiyang2022,luyaojie,hongyu,xianpei,sunle}@iscas.ac.cn {jcaoap, scc}@cse.ust.hk mwenaa@hust.edu.cn benhe@ucas.edu.cn
footnotetext: *These authors contributed equally to this work.
1 Introduction
@@
Large language models (LLMs) have shown advanced proficiency in various domains, including code generation (Liu et al., 2024; Du et al., 2024), defect detection (Yang et al., 2024a) and program repair (Xia and Zhang, 2023; Zhong et al., 2024; Hu et al., 2024). Benchmarks such as HumanEval (Chen et al., 2021) and SWE-bench (Jimenez et al., 2023) were introduced to measure LLMs’ capabilities, providing insights into LLMs’ strengths and weaknesses.
@@
Recent studies (Cao et al., 2024a; Chai et al., 2024; Chen et al., 2024) have spotted two significant biases in current benchmarks. First, Programming language bias. As pointed out by prior studies (Cao et al., 2024a; Chai et al., 2024; Chen et al., 2021; Austin et al., 2021), Python dominates code generation benchmarks with over 95% involvement. Other programming languages (PLs) such as Java and C/C++, despite their popularity and availability, gain less exploration. Second, Coding task bias. Most coding benchmarks focus on code generation tasks (i.e., giving descriptions in natural language and generating the program, as shown in Figure 1 (A)), while code reasoning (i.e., given the program, reasoning the input or output of the program, as shown in Figure 1 (B - D)) is seldom evaluated (Chen et al., 2024). A recent work introduced a code reasoning benchmark (Gu et al., 2024; Chen et al., 2024), while it is only in Python. Figure 1 (C - D) shows that simply changing PLs from Python to C++ can turn a correct reasoning into an incorrect one. More discussion about this task can be found in Appendx J.
@@
However, constructing multi-lingual benchmarks is not a trivial task. First, human annotation can be expensive. As reported by recent work (Chai et al., 2024), they spent a total of \$12,000 US dollars for human annotators, providing the working environment, free meals, souvenirs, and free GPT-4 interface usage to construct their multi-lingual benchmark. Second, automated translation does not perform well. The latest studies (Yin et al., 2024) show that even the best LLM (i.e., ChatGPT) can only achieve an average of 64% success translation rate, which is far from practice. Rule-based translation (Cassano et al., 2023; Ling et al., 2022) usually suffers from generalizability issues, making them limited in handling prescribed code structures. Additionally, multi-lingual solutions from contest websites such as LeetCode and Codeforces were included in most LLMs training sources, thus suffering from data contamination issues (Cao et al., 2024b).
@@
To fill the above research gaps, we introduce CruxEval-X, a multi-lingual code reasoning benchmark that contains 19 popular PLs, including C++, Rust, Java, etc. , expanded from CruxEval (Gu et al., 2024), a code reasoning benchmark written in Python. For each PL in CruxEval-X, there are at least 600+ functions. In total, there are 12,660 subjects along with 19K test cases for input/output reasoning.
@@
Noteworthy that the pipeline of constructing CruxEval-X works in a fully automated manner. It first translates the test cases by transition rules adapted from prior work (Cassano et al., 2023), then iterates the generation-and-repair process intensively. In particular, the transition rules are formulated to cross the language barriers. For example, Python employs a dynamically typed system where types are determined at runtime, whereas C++ uses a statically typed system requiring explicit type declarations at compile time. The rules facilitate the translation of the test cases. Additionally, inspired by prior work (Yin et al., 2024; Rozière et al., 2022), we employ a test-guided manner (Rozière et al., 2022) to generate the translation and iteratively repair the generated code using execution feedback (e.g., compilation error, runtime error) (Yin et al., 2024).
@@
Through extensive experiments on 24 mainstream LLMs, we observe several interesting findings. First, in multiple PLs, the input reasoning and output reasoning capabilities of LLMs are comparable. Also, there is a noticeable correlation between certain PLs (e.g., JavaScript and TypeScript show a positive correlation, while Racket consistently yields the worst results). More interestingly, we observe that even if a model is only trained on Python (e.g., phi-1 and phi-1.5), it still can reach a 16% $\sim$ 26% output reasoning success rate in other PLs, compared with 25.6% in Python. The finding indicates the cross-language generalization of LLMs.
@@
The contributions can be summarized as follows: (1) We introduce CruxEval-X, a multi-lingual code reasoning benchmark that contains 19 popular PLs. (2) We introduce an automated code translation pipeline that adopts a test-guided and iterative generate-and-repair practice. (3) We evaluate 20+ LLMs and yield inspiring findings.
@@
2 Benchmark Construction
In this section, we detailed the construction process of CruxEval-X in Figure 2. It can be divided into three main steps. First, we translate function signatures via mapping variable type annotations (Step 1 in Figure 2). Then we employ a rule-based approach to translate Python test cases into other PLs (Step 2 in Figure 2). Finally, we integrate multiple LLMs to translate code by iterating the generation-and-repair process (Step 3 in Figure 2).
2.1 Step I. Function Signature Translation
@@ 2 Benchmark Construction
Then, we adopt the rules as prior work (Cassano et al., 2023) to map the types from Python to other PLs. In particular, we identify the data types in the annotated Python signature (e.g., parameter types, return types), mapping the types from Python to other PLs according to the rules, then structuring the signature in the corresponding PLs. Take the example in Figure 2 Step I, the Python signature def f(s1:str, s2:str) -> str is translated into that in C++ (std::string f(std::string s1, std::string s2). After translating the tests, all 800 subjects in Python can be translated, as shown in Table 1.
2.2 Step II. Test Suites Translation
@@
We employ a test-guided approach to ensure the correctness of the translation results, which necessitates test cases in various PLs. Prior works (Athiwaratkun et al., 2022; Cassano et al., 2023) provided various rules for mapping Python test cases to other PLs. We adopt the mapping rules from MutiPL-E (Cassano et al., 2023) to assist the transition of our test suites.
@@
However, their rules have limited support for type handling (e.g., they cannot handle a list with hybrid types), Thus, to maximize the success rate, we made two improvements to enhance the rules. First, we enhance handling structured types such as List and Dict. For example, when handling C#, we add an equality function to check whether two Dict types are equal. Second, when dealing with variables that have complex types that are not as well-supported in some other PLs, we transform these variables into more generic types without significantly altering the original function’s functionality. For example, we change type List[Union(str, int)] into List(str) if the function keeps the same functionality. A small portion of the data that cannot be converted is discarded. The result of Step II is shown in Table 1. Further details can be found in the Appendix B.
2.3 Step III. Iterative Generation & Repair
After translating tests, dependencies, and signatures, we employ multiple LLMs to iteratively translate Python code into target PLs. Each LLM performs two steps: generation and repair. Problems that pass all test cases are excluded from subsequent iterations. Below, we will elaborate on these two substeps in detail.
2.3.1 Generation.
Relying on a single LLM’s limited generations often results in a low translation success rate (Yin et al., 2024). To address this, we propose a multi-round generation method that interacts with the testing environment to determine whether to continue iterating.
@@
Let $A_{0}$ denote the number of correctly translated codes, and $U$ the total number of problems. For unresolved problems, LLM $M$ generates results over multiple rounds. Let $A_{i}$ represent the number of correct results in the $i$-th round. The maximum rounds is $N$. Early stopping occurs if the increase in correct results compared to $k$ rounds prior, $A_{i-k}$, falls below a threshold $\delta$. The formula is as follows:
@@
Here, $O_{i}$ is the $i$-th round code generated from $P(O_{i}\ |\ U-A_{i-1};M)$, and $Correct(.)$ calculate the correct results. To leverage diverse LLMs’ strengths, we use GPT-3.5-Turbo for initial generation (see Table 1, “w/o Iter”) and DeepseekCoder-33B-Instruct for further generation. This pipeline improves upon single-LLM generation, as demonstrated in our final results. (see Table 1, “w/ Iter”)
2.3.2 Repair
@@
Simply generating code will still result in many errors that LLMs cannot solve. Therefore, after the generation step of each LLM, we provide them with error messages for error correction. Multiple iterative error correction is costly with limited benefits. (Chen et al., 2023), so we only perform error correction once after each LLMs.
@@
After GPT3.5-Turbo generation, we directly provide the LLM with the erroneous code along with the error messages for correction. After the Iterating Generation of DeepseekCoder-33B-Instruct, since the untranslated code can produce numerous incorrect code snippets after multiple rounds of iteration, which may contain the same errors. We first use simhash to deduplicate the erroneous code and then proceed with error correction on the deduplicated code. The error correction in each phase uses the LLM used in that phase. with a temperature setting of 0.
2.3.3 Multiturn Repair Based on Overlap
@@
After completing the steps above, we first calculate the intersection of correctly answered questions in different PLs. To our surprise, there were only 333 questions that all PLs answered correctly. However, 563 problems have been successfully translated correctly by at least 16 PLs. Upon analyzing the questions our LLM failed to solve, we find that each PL has its difficulties in translating from Python. For example, in Julia, the index for arrays and other collection types starts from 1, which differs from Python. The details of the difficulties can be found in Appendix C.
@@
Based on these observations, we conduct final generation and iterative error correction on the questions that are correctly translated by more than 15 PLs. Due to the small amount of data, we utilized GPT-4o for generation and error correction. During the generation process, we provide the LLM with three corresponding typical examples based on the difficulties we find. The overlap is increased to 462 after repair of GPT-4o.
@@
We manually refined 38 questions that GPT-4 nearly solved, expanding our dataset to 500 entries. We determined that 500 entries are sufficient to distinguish the effectiveness of the LLMs. Therefore, we use these 500 entries as our CruxEval-X benchmark. The final result of our pipeline is shown in Table 1 under the column “w/ Iter”. The experimental setups and prompts can be found in Appendix D, and detail statics of our benchmark is shown in Appendix I.
2.4 Quality Analysis
After constructing the dataset, we evaluate the benchmark’s quality from four perspectives. First, Accuracy: We use a test-guided approach to ensure that each data point passes test cases, guaranteeing correctness. Second, Reliability: We build on the high-quality CRUXEVAL dataset, ensuring problems avoid randomness or multi-threading, enhancing reliability. Third, Data Leakage: We compare with Stack v2 (67.5TB of GitHub data), finding only 0.8% overlap, indicating minimal leakage risk. Fourth, Data Bias: We assess potential bias by generating evaluation data with other LLMs using the same method as described above. Results show no significant advantage for any LLM on its own generated data. The possibel reason is the distinct tasks in generation and evaluation phases. Detailed results and analysis are in Appendix A.
@@ 2.4 Quality Analysis
3 Experiments
3.1 Experiment Setup
Input Reasoning Performence
Models Size cs cpp d go java js jl lua pl php py r rkt rb rs scala sh swift ts
GPT-4o - 70.2 64.6 71.6 75.4 69.8 73.2 67.0 73.0 70.2 74.8 70.6 74.4 67.4 72.0 73.6 65.4 70.6 74.2 74.0
GPT-4o-mini - 58.8 52.2 60.6 62.0 57.2 59.6 56.2 63.4 57.4 61.0 59.6 60.4 51.2 61.6 61.2 52.6 57.2 63.4 61.2
GPT-3.5 Turbo - 52.2 39.2 50.2 53.4 55.4 50.0 47.0 53.2 47.6 52.2 51.6 48.6 45.4 49.6 53.0 54.2 47.6 58.2 48.4
Deepseekcoder-V2 236B 63.8 57.0 66.6 64.0 64.8 67.0 58.4 62.0 61.4 64.2 64.0 65.8 58.0 63.2 63.6 58.2 62.4 62.6 66.6
Qwen2-Instruct 72B 52.0 54.2 49.6 55.4 50.0 51.6 51.0 51.2 47.8 55.2 52.4 53.2 47.8 54.4 57.2 50.6 52.4 51.6 52.0
CodeLlama-Python 34B 38.8 40.0 39.2 39.0 41.4 45.8 44.8 45.0 43.2 48.0 46.8 42.2 38.8 44.0 44.2 43.0 44.6 45.0 44.0
CodeLlama-Instruct 34B 44.6 48.4 43.8 46.0 44.4 52.6 50.4 49.4 46.0 52.0 51.2 48.4 42.4 48.2 48.6 48.0 46.2 49.4 53.2
CodeLlama 34B 40.4 44.6 45.6 41.2 39.0 50.0 49.0 47.0 46.6 48.8 49.8 47.6 39.8 46.6 46.8 44.6 44.4 50.0 48.6
WizardCoder-V1.1 33B 44.8 25.4 46.4 47.6 48.4 45.6 49.2 48.8 44.6 50.0 50.0 45.0 42.4 49.2 48.2 48.2 45.4 51.0 46.4
Deepseekcoder-instruct 33B 46.0 43.6 49.8 49.0 46.8 48.8 47.0 50.0 46.8 52.0 51.8 48.2 41.6 52.0 48.4 47.0 48.2 52.2 49.6
Deepseekcoder-base 33B 41.2 42.8 43.2 45.6 43.8 46.0 47.6 47.4 47.2 48.6 49.2 50.6 42.8 47.4 46.8 44.0 46.4 48.2 45.0
Starcoder2 15B 41.4 43.8 51.6 45.2 42.6 44.0 48.2 44.6 44.8 49.8 46.6 45.8 45.0 49.0 46.6 37.0 47.4 52.2 46.2
WizardCoder-V1.0 15B 29.2 30.0 30.6 28.6 29.6 33.0 34.8 33.6 36.2 36.8 33.2 33.4 36.4 33.6 33.0 29.0 35.0 34.0 32.4
Starcoder 15B 28.2 30.0 33.0 33.2 33.4 35.2 34.4 31.6 34.0 36.4 34.8 33.4 36.6 35.0 34.8 27.4 37.0 30.8 33.2
phi-3-instruct 14B 31.8 26.0 38.8 36.4 37.2 42.4 36.2 37.2 35.6 41.2 43.4 39.2 24.4 36.0 36.8 38.0 33.6 41.2 42.8
Llama-3-Instruct 8B 37.0 36.4 35.0 38.6 36.2 38.4 39.6 40.0 36.2 36.6 38.4 42.2 24.2 35.8 37.6 38.0 31.6 42.2 38.2
CodeQwen1.5-Chat 7B 42.8 42.0 43.0 46.4 44.6 43.8 42.2 42.8 41.6 44.8 43.0 43.4 38.2 43.6 42.0 39.4 46.6 45.8 43.6
CodeLlama-Instruct 7B 38.6 36.0 38.4 38.4 38.2 39.6 42.2 43.4 36.4 40.4 41.0 41.0 38.8 41.6 37.6 42.6 39.6 40.2 41.0
CodeLlama-hf 7B 36.4 36.2 36.8 34.6 36.4 36.6 40.2 39.6 36.0 39.4 40.2 40.0 36.6 39.2 35.4 37.8 36.8 39.2 38.8
Deepseekcoder-instruct 6.7B 35.0 37.0 35.6 40.4 35.0 36.6 39.2 38.8 39.4 42.2 38.2 42.0 37.2 40.2 37.4 36.8 42.8 40.8 34.2
Deepseekcoder-base 6.7B 38.8 42.4 41.2 43.2 40.4 43.6 42.6 42.8 41.6 46.4 41.4 46.2 43.0 44.6 41.6 40.8 44.8 43.4 41.8
CodeGen-multi 6B 28.8 25.4 6.2 25.6 36.2 25.2 17.4 24.4 38.4 22.8 22.6 27.2 16.2 6.4 18.8 31.0 48.6 32.4 25.2
phi-1_5 1.3B 29.2 16.0 13.2 25.8 26.8 9.8 30.4 26.6 17.8 26.6 25.8 8.4 6.6 1.4 25.2 30.4 34.4 26.6 30.8
phi-1 1.3B 0.2 7.0 9.6 3.6 2.8 17.0 19.0 17.4 23.6 9.2 11.8 9.4 11.2 6.8 5.4 1.8 19.8 14.0 14.0
Average 40.4 38.3 40.8 42.4 41.7 43.1 43.1 43.9 42.5 45.0 44.1 43.2 38.0 41.7 42.7 41.1 44.3 45.4 43.8
Output Reasoning Performence
Models Size cs cpp d go java js jl lua pl php py r rkt rb rs scala sh swift ts
GPT-4o - 75.0 74.8 71.4 77.0 73.2 77.6 73.6 74.8 74.0 75.4 75.4 72.0 70.8 74.0 74.4 71.8 71.6 76.0 76.4
GPT-4o-mini - 63.0 63.0 61.4 63.4 54.0 61.8 57.8 60.0 57.4 64.2 61.6 59.6 56.6 61.2 61.8 61.2 56.2 63.0 61.2
GPT-3.5 Turbo - 54.2 43.2 56.0 53.2 43.6 56.2 54.2 54.6 51.8 55.2 57.2 49.4 48.0 56.4 54.6 56.4 51.0 57.8 53.6
Deepseekcoder-V2 236B 66.6 66.2 63.4 68.0 67.6 65.4 64.8 63.6 63.0 67.4 66.8 63.0 62.2 65.2 65.8 63.2 58.8 67.8 66.4
Qwen2-Instruct 72B 51.2 50.2 51.6 53.6 38.2 52.0 51.0 49.0 45.8 50.8 51.2 45.0 46.8 50.8 51.0 51.0 45.6 50.4 53.2
CodeLlama-Python 34B 41.4 44.8 45.6 41.8 41.4 45.4 45.2 42.8 43.6 43.8 43.8 42.4 38.6 42.8 46.6 43.8 42.0 44.4 44.8
CodeLlama-Instruct 34B 44.4 46.2 45.8 46.8 40.6 47.4 45.6 42.8 44.0 44.8 44.0 40.2 38.2 44.2 46.4 43.8 40.6 45.2 45.0
CodeLlama 34B 44.6 47.8 44.2 45.2 38.4 47.0 45.8 42.8 43.8 46.4 46.4 38.8 38.4 45.4 47.2 47.4 43.8 47.6 47.4
WizardCoder-V1.1 33B 47.0 46.8 45.8 44.2 50.8 50.0 47.0 46.0 45.2 51.4 49.6 44.0 42.4 48.2 47.8 45.0 44.4 48.0 49.8
Deepseekcoder-instruct 33B 52.0 51.4 49.0 48.8 53.2 55.0 50.4 50.4 50.0 53.0 52.2 48.2 46.6 52.8 50.6 48.0 49.4 52.8 53.6
DeepseekCoder-base 33B 48.2 50.0 46.0 48.6 49.2 51.4 46.8 48.0 48.4 52.0 49.8 45.2 46.4 49.0 46.2 47.6 46.0 49.2 51.2
Starcoder2 15B 46.0 47.4 47.2 49.0 48.4 50.0 49.2 44.8 49.4 48.4 48.4 47.2 45.0 51.0 48.8 45.2 45.8 49.6 48.6
WizardCoder-V1.0 15B 25.2 30.0 30.6 33.2 26.8 33.6 30.2 30.2 31.0 33.0 34.0 31.6 29.6 32.8 31.2 31.2 29.8 34.2 34.0
Starcoder 15B 20.4 31.6 31.8 31.0 18.4 33.4 32.2 31.8 29.8 32.6 32.6 30.0 29.2 33.4 32.6 30.0 30.2 33.0 33.0
phi-3-instruct 14B 34.2 37.6 39.0 31.0 34.2 41.6 41.2 34.4 35.8 37.8 42.4 36.6 24.6 42.2 37.4 36.2 37.2 41.4 43.0
Llama-3-Instruct 8B 32.0 30.8 31.2 31.4 25.0 35.0 31.4 34.0 29.6 27.0 33.6 27.2 28.0 31.8 34.4 33.8 32.0 36.4 33.8
CodeQwen1.5-Chat 7B 37.8 40.2 40.2 40.6 35.4 43.6 42.6 40.4 39.6 43.0 41.4 38.2 39.0 44.6 42.0 35.0 38.2 43.8 42.2
CodeLlama-Instruct 7B 32.2 35.6 34.4 35.0 24.4 38.2 35.2 32.2 34.2 36.0 35.4 32.0 29.6 37.0 37.4 33.0 33.0 34.6 38.8
CodeLlama-hf 7B 32.6 34.4 33.8 33.4 28.4 38.0 35.2 34.4 35.2 38.0 34.4 32.6 30.8 34.8 36.8 33.4 31.0 35.0 38.2
Deepseekcoder-instruct 6.7B 34.8 41.8 40.4 39.4 32.8 47.6 42.6 38.8 42.0 43.8 43.6 40.8 39.2 43.2 41.8 40.6 37.8 43.2 44.0
Deepseekcoder-base 6.7B 41.2 46.2 43.2 42.8 42.6 44.8 46.0 41.0 40.4 41.8 44.8 42.8 43.0 42.6 42.0 43.2 40.6 47.6 45.4
CodeGen-multi 6B 21.4 23.6 25.0 26.4 21.6 22.8 22.8 23.8 20.4 25.2 24.8 23.4 17.8 24.0 25.2 22.0 22.2 25.0 21.4
phi-1_5 1.3B 16.0 26.0 24.8 22.6 15.8 23.0 23.6 21.2 22.0 22.2 25.6 21.8 16.8 19.6 22.0 21.6 17.6 25.6 25.2
phi-1 1.3B 5.8 9.0 13.2 14.8 4.6 20.8 19.2 15.8 15.6 18.6 22.4 17.6 10.4 18.0 16.4 11.0 16.4 19.2 19.0
Average 40.3 42.4 42.3 42.6 37.9 45.1 43.1 41.6 41.3 43.8 44.2 40.4 38.3 43.5 43.4 41.5 40.1 44.6 44.6